A Candy Barrel shop manager mixes M%M's worth $2.00 per pound with trail mix worth 1.50 per pound. Find how many pounds of each she would use to get 50 pounds of party mix worth $1.80.
Answer by Paul(988)

2x+1.50(50-x)=50(1.8)
2x+75-1.5x=90
0.5x=15
x=30

50-30=20

Hence, for $2.00 about 30pounds of mixture was added, and for $1.5 about 20 pounds of mixture was added.
Paul.